# When predict can also explain: few-shot prediction to select better neural latents

### v3

@@ Appendix F Analysis of SOTA models
Appendix G Validating cross-decoding column-mean as a proxy of ground truth distance in HMMs
For SOTA models, we don’t have ground truth and therefore use cross-decoding as a proxy. We validate this approach in the HMM setting, where we can compute cross-decoding among student models, while also having access to ground truth, i.e., the teacher. As Fig. 9 shows, the novel cross-decoding metric is highly correlated to the ground truth metric of interest $\mathcal{D}_{\text{T}\rightarrow\text{S}}$.
Figure 9: For HMM students with high co-smoothing $\mathcal{Q}_{\text{S}}>\mathcal{Q}_{\text{T}}-10^{-3}$ (and therefore low $\mathcal{D}_{\text{S}\rightarrow\text{T}}$ 1D), the cross-decoding metric $\langle\mathcal{D}_{u\rightarrow v}\rangle_{u\in\text{students}}$ is correlated to ground truth distance $\mathcal{D}_{\text{T}\rightarrow\text{S}}$ and uncorrelated to $\mathcal{D}_{\text{S}\rightarrow\text{T}}$.
Next, in Fig. 10 we replicate the comparison in Fig. 6 of the main text, with the HMMs instead of SOTA models. Despite very different LVM architectures and very different datasets (synthetic versus real neural data), the results are strikingly similar.
Taken together, these results reinforce our use of the novel cross-decoding metric as a proxy to $\mathcal{D}_{\text{T}\rightarrow\text{S}}$ for SOTA models on real data where there is no access to ground truth, i.e., no teacher model T.
Figure 10: Few-shot co-smoothing validated with cross-decoding in HMMs: a repeat of main text Fig. 6, now in the HMM setting. For HMMs, $v\in\text{students}$, with near-optimal co-smoothing, $\mathcal{Q}_{v}>\mathcal{Q}_{\text{T}}-10^{-3}$, few-shot co-smoothing scores $\langle\mathcal{Q}^{k}_{v}\rangle$, with $k=6$, negatively correlate with the cross-decoding metric $\langle\mathcal{D}_{u\rightarrow v}\rangle_{u}$, used as proxy for the distance from ground truth metric $\mathcal{D}_{\text{T}\rightarrow\text{S}}$. Meanwhile, co-smoothing scores $\mathcal{Q}$ are uncorrelated with the same.
Appendix H How to choose $k$ and $s$?
We define $\mathcal{Q}^{k}$ the $k$-shot co-smoothing score: the co-smoothing score given by predictions from decoder $g^{\prime}$ trained with only $k$ trials of the $k$-out neurons and the corresponding latents given by the encoder $f$ (section 5 and Fig. 3). As this can be variable across random $k$-trial subsets we report the average $k$-shot co-smoothing, $\langle\mathcal{Q}^{k}\rangle$, averaging over $s$ decoders each independently trained on random resamples of $k$-trials. Here we report how the results change with $k$, offering guidelines on how to choose $k$ and $s$.
We first analyse the student-teacher HMMs from 4. In Fig. 11 we show several quantities as a function of $k$. We see that small $k$ maximimally separates two extreme models and the scores converge for $k\to\infty$. However at small $k$, scores $\mathcal{Q}^{k}$ from single models are also more variable, therefore more resamples $s$ are required for a good estimate of the mean $\langle\mathcal{Q}^{k}\rangle$. We choose $s:=\lfloor\frac{S_{\text{train}}}{k}\rfloor$ and find that $k\approx 6$ gives us the best correlation to ground-truth measure (Fig. 11 bottom-right).
We do a similar analysis for LFADS models on the mc_maze_20 dataset. In Fig. 12 we show several values of $Q^{k}$ (appendix F) for several random samples of $k$-trials, and at various values of $k$. We find that for $k$ values including and below $k=32$, scores are negative, and at $k=4$ scores are even worse and vary by orders of magnitude. Among the values we checked, we found $k=128$ to be the smallest value with positive and low-variance $\mathcal{Q}^{k}$. Thus, in Fig. 6 we use an intermediate value of $k=128$ and $s:=\lfloor\frac{S_{\text{train}}}{k}\rfloor$.
Figure 11: Choosing $k$ and $s$: analysis with HMMs. Top-left: Average $k$-shot co-smoothing as a function of $k$ for three models, the teacher T, a good and a bad student as (see 2). Top-right Standard deviation of $k$-shot co-smoothing values across resamples. Bottom-left Signal to noise ratio, ratio of standard deviation of $k$-shot co-smoothing values across models vs with models. Bottom-right: Pearson’s correlation of average $k$-shot score and the ground-truth decoding measure, for models with high co-smoothing $Q$, as reported in Fig. 4B for $k=6$. Here we take $s:=\lfloor\frac{S_{\text{train}}}{k}\rfloor$.
Figure 12: $k$-shot scores (without averaging over $s$ resamples) for LFADS models on the mc_maze_20 dataset, as a function of $k$. $\mathcal{Q}^{k}=0$ is a baseline score obtained by reporting the mean firing rate for each neuron. For small $k$ scores fall below 00 and become highly variable.
Appendix I Dimensions of datasets
@@
We analyse three datasets in this work. Two synthetic datasets generated by an HMM (ground truth in Fig. 2), an LGSMM (appendix E) and the mc_maze_20 dataset from the Neural Latent Benchmarks (NLB) suite (Pei et al., 2021; Churchland et al., 2010). In table 1, we summarise the dimensions of these datsets. To evaluate $k$-shot on the existing SOTA methods while maintaining the NLB evaluations, we conserved the forward-prediction aspect. During model training, models output rate predictions for $T^{\text{fp}}$ future time bins in each trial, i.e., equation 1 and equation 2 are evaluated for $1\leq t\leq T^{\text{fp}}$ while input remains as ${\bm{X}}_{1:T,\text{in}}$. Although we do not discuss the forward-prediction metric in our work, we note that the SOTA models receive gradients from this portion of the data.
@@
In mc_maze_20 we reuse held-out neurons as $k$-out neurons. We do this to preserve NLB evaluation metrics on the SOTA models, as opposed to re-partitioning the dataset resulting in different scores from previous works. This way existing co-smoothing scores are preserved and $k$-shot co-smoothing scores can be directly compared to the original co-smoothing scores. The downside is that we are not testing the few-shot on ‘novel’ neurons. Our numerical results (Fig. 6) show that our concept still applies.
@@
Table 1: Dimensions of real and synthetic datasets. Number of train and test trials $S^{\text{train}}$, $S^{\text{test}}$, time-bins per trial for co-smoothing $T$, and forward-prediction $T^{\text{fp}}$, held-in, held-out and $k$-out neurons $N^{\text{in}}$, $N^{\text{out}}$, $N^{\text{$k$-out}}$.
@@
Synthetic HMM $2000$ $100$ $10$ $20$ $50$ $50$
Synthetic LGSSM (appendix E) $20$ $500$ $10$ $5$ $30$ $30$
NLB mc_maze_20 (Pei et al., 2021; Churchland et al., 2010) $1721$ $574$ $35$ $10$ $127$ $55$ 55555555222In mc_maze_20 we use the same set of neurons for $N^{\text{out}}$ and $N^{\text{$k$-out}}$.
Appendix J Code repositories
The experiments done in this work are largely based on code repositories from previous works. The code developed here is in https://osf.io/4bckn/?view_only=73b3aee9a8eb43e8bb3b286c800c6448. Table 2 provides links to the code repositories used or developed in this work.
Table 2: Summary of key repositories used in this paper
Repository Forked from Citations
anonymous repo https://github.com/neurallatents/nlb_tools (Pei et al., 2021)
anonymous repo https://github.com/trungle93/STNDT (Le & Shlizerman, 2022; Ye & Pandarinath, 2021; Pei et al., 2021; Nguyen & Salazar, 2019; Huang et al., 2020)
anonymous repo https://github.com/arsedler9/lfads-torch (Sedler & Pandarinath, 2023; Pandarinath et al., 2018; Keshtkaran et al., 2022)
Appendix K Compatibility and consistency of cross-decoding across LVM architectures
In this section we analyse the cross-decoding approach, pooling together the SOTA models from the two architectures: STNDT and LFADS, all trained on the same mc_maze_20 dataset. We filtered models to those with near-SOTA co-smoothing, specifically $0.360.348<\mathcal{Q}<$ 0.36, resulting in $75$ LFADS and $40$ STNDT models. Note that this included LFADS models which were not in the main text Fig. 6.
Fig. 13 shows the cross-decoding matrix $\mathcal{D}_{u\rightarrow v}$ for all pairs of models $(u,v)$ in this combined set, as computed in section 7 and appendix F. The cross-decoding matrix reveals a block structure, suggesting larger decoding errors for model pairs from different architectures versus model pairs within the same architecture. Crucially, on top of this block structure, we see clear continuation of columns. This implies that models that are extraneous in one class are also judged as extraneous by the other class. This is summarised in Fig. 14, where we compare column means for each model over the ‘same architecture pool’ and ‘other architecture pool’. Thus, cross decoding can be used across architectures. One should note, however, that having an unbalanced sample from the two classes could bias scores to be lower for the larger class. Finally, we use the combined cross-decoding matrix to repeat the analysis of the main text, but combining both model types. Fig. 15 shows that our conclusions hold – co-smoothing is uncorrelated with cross-decoding, while few-shot is correlated.
Figure 13: Cross decoding matrix $\mathcal{D}_{u\rightarrow v}$ for all model pairs $(u,v)$ from the combined set of $75$ LFADS and $40$ STNDT models on mc_maze_20 with near-SOTA co-smoothing $0.360.345<\mathcal{Q}<$ 0.36. The colourmap saturates at the upper 99% quantile of scores in the matrix to better visualise the bulk of the data.
Figure 14: Architecture-wise column means of the cross-decoding matrix. Compare their cross-decoding column means for input models from models of the same architecture versus models of a different architectures, i.e., $\langle\mathcal{D}_{u\rightarrow v}\rangle_{u\in U\backslash\{v\}}$ for target models $v\in V$. LEFT: $V$ is the population of STNDT models and RIGHT: LFADS. All models in $U\cup V$ have near-SOTA co-smoothing, in the range $0.360.345<\mathcal{Q}<$ 0.36. We do not include the self-decoding scores $\mathcal{D}_{u\rightarrow u}$ as these are trivially near-zero and bias the results.
Figure 15: Few-shot scores correlate with the proxy of distance to the ground truth, even in a mixed population of architectures. We repeat the analysis in main text Fig. 6, pooling together the STNDT and LFADS models with high co-smoothing scores $\mathcal{Q}>0.348$ and compute cross-decoding $\langle\mathcal{D}_{u\rightarrow v}\rangle_{u}$ for the combined population. To replicate the main text result, we plot models in narrow $\mathcal{Q}$ range, i.e., an upper co-smoothing limit of $\mathcal{Q}<0.355$, while ensuring that models of both architectures are included.
Appendix L Theoretical analysis of ridgeless least squares regression with extraneous noise
$x_{i}=z^{*}_{i}+\epsilon_{i},$ (24)
${\bm{z}}_{i}:=\begin{bmatrix}z^{*}_{i}&\bm{\xi}_{i}\end{bmatrix}^{T},$ (25)
$\hat{\bm{w}}=\operatorname*{arg\,min}\left\{\|w\|_{2}:w\text{ minimizes }\sum_{i=1}^{k}\|x_{i}-{\bm{w}}^{T}{\bm{z}}_{i}\|^{2}_{2}\right\}.$ (26)
Note that, by construction, the true mapping is:
${\bm{w}}^{*}=\begin{bmatrix}1&0&\dots&0\end{bmatrix}^{T}.$ (27)
Test loss or risk is a mean squared error:
given a test sample ${\bm{z}}_{0}$. The error can be decomposed as:
The scenario described above is a special case of (Hastie et al., 2022). What follows is a direct application of their theory, which studies the risk $R$, in the limit $k,p\to\infty$ such that $p/k\to\gamma\in(0,\infty)$, to our setting. The alignment of the theory with numerical simulations is demonstrated in Fig. 16.
Claim 1.
$\gamma<1$, i.e., the underparameterised case $k>p$.
$B=0$ and the risk is just variance and is given by:
with no dependence on $\sigma_{\text{ext}}$.
Proof.
This is a direct restatement of Proposition 2 in (Hastie et al., 2022). ∎
Claim 2.
$\gamma>1$, i.e., the overparameterised case $k<p$.
The following is true as $k,p\to\infty\text{ and }p/k\to\gamma$:
Proof.
For the non-isotropic case (Hastie et al., 2022) define the following distributions based on the eigendecomposition of $\Sigma$.
$d\widehat{G}(s)$ $=\delta(s-1)$ (34)
In the limit $p\to\infty$ we take $d\widehat{H}(s)\approx\delta(s-\sigma_{\text{ext}}^{2})$. This greatly simplifies calculations and nevertheless provide a good fit for numerical results with finite $k$ and $p$. We solve for $c_{0}(\gamma,\widehat{H})$ using equation 12 in (Hastie et al., 2022).
$\gamma c_{0}=\frac{1}{(\gamma-1)\sigma_{\text{ext}}^{2}}$ (35)
We then compute the limiting values of $B$ and $V$:
$V$ $=\sigma_{\text{obs}}^{2}\gamma c_{0}\sigma_{ext}^{2}.$ (37)
Substituting $\gamma c_{0}$ completes the proof. ∎
Comments:
  • The extraneous noise, $\sigma_{\text{ext}}$, only affects the risk of ridgeless regression for $k<p$,
  • moreover, it only affects the bias term not the variance.
  • Observation noise only affects the variance term, therefore the dependence of the risk on $\sigma_{\text{ext}}$ would occur even if $\sigma_{\text{obs}}=0$, i.e., in the absence of observation noise.
Figure 16: Theoretical analysis of few-shot ridgeless least squares regression for $p=21$ and $p=51$. Dots are from numerically solved regressions for finite $k$ and $p$, and dashed lines are the theoretical predictions from the claims 1 and 2, based on (Hastie et al., 2022).
Appendix M Theoretical analysis of prototype learning for binary classification with extraneous noise
Teacher latents are distributed as $p(z_{i}^{*})=\frac{1}{2}\delta(z_{i}^{*}-\frac{1}{\sqrt{2}})+\frac{1}{2}\delta(z_{i}^{*}+\frac{1}{\sqrt{2}})$, that is either $\frac{1}{\sqrt{2}}$ or $-\frac{1}{\sqrt{2}}$ with probability $\frac{1}{2}$, representing two classes $a$ and $b$ respectively.
[…]
In our case this reduces to:
$SNR$ $\approx\frac{\sqrt{Mk}}{\sigma_{\text{ext}}^{2}}$ (40)
Comments:
  • Classification error monotonously decreases with number of samples $k$, $\epsilon\to 0$ as $k\to\infty$, for all finite values of $\sigma_{\text{ext}}$.
  • Classification error monotonously increases with extraneous noise $\sigma_{\text{ext}}$, significantly deviating from zero when $\sigma_{\text{ext}}^{2}\approx\sqrt{Mk}$.
Figure 17: Theoretical analysis of few-shot binary classification with prototype learning with $M=20$. Dots are numerically computed errors and dashed lines are the theoretical predictions from equation 40, based on (Sorscher et al., 2022).
Appendix N Theoretical analysis of few shot learning in HMMs
@@
Consider a student-teacher scenario as in section 4. We let $T=2$ and use a stationary teacher $z^{(i)}_{1}=z^{(i)}_{2}$. Now consider two examples of inferred students. To ensure a fair comparison, we use two latent states for both students. In the good student, $\xi$, these two states statistically do not depend on time, and therefore it does not have extraneous dynamics. In contrast, the bad student, $\mu$, uses one state for the first time step, and the other for the second time step. A particular example of such students is:
@@
$\begin{array}[]{ccc}\hat{B}_{1}(\xi)=\frac{0.5(C_{1}+C_{2})}{0.5kT}&&\hat{B}_{1}(\mu)=\frac{C_{1}}{k}\\ \hat{B}_{2}(\xi)=\frac{0.5(C_{1}+C_{2})}{0.5kT}&&\hat{B}_{2}(\mu)=\frac{C_{2}}{k}\\ \end{array}$ (45)
@@
Figure 18: Test loglikelihood, $k$-shot binomial regression for two, $2$-state model latents: a compact latent $\xi$ and an extraneous latent $\mu$. Dots are numerical and dashed lines are theory.
Appendix O Cross decoding matrices and example latents for LFADS AND STNDT
Figure 19: Cross-decoding as a proxy for distance to the ground truth in near-SOTA models. 200 LFADS models (left) and 120 STNDT models (right) were trained on the mc_maze_20 dataset then selected for high-cosmoothing (appendix F). The latents of each pair of models were decoded from one another, and the decoding error is shown in the matrices. Good models are expected to be decoded from all other models, and hence have low values in their corresponding columns. Bottom left: Trajectories of two LFADS models, with the lowest (left and in a green box, best model$:=\text{argmin}_{v}\langle\mathcal{D}_{u\rightarrow v}\rangle_{u}$) and highest (right and in the red box, the worst model$:=\text{argmax}_{v}\langle\mathcal{D}_{u\rightarrow v}\rangle_{u}$) column averaged cross-decoding errors, projected onto their leading two principal components. Scores for these models are indicated in Fig. 6 by the arrows. Each trace is the trajectory for a single trial, starting at a green dot and ending at a red dot. Bottom right Same for STNDT.

### v4

@@ Methods
Cycle consistency
We evaluate cycle-consistency [8, 29] for a model $u$ also using a linear mapping from its rate predictions $\bm{R}$} back to its latents $\hat{\bm{Z}}$ implemented with sklearn.linear_model.LinearRegression:
$\left(\hat{\bm{Z}}^{(i)}_{t,:}\right)_{u}=h_{\bm{r}\rightarrow\bm{\hat{z}}}\left(\left({\bm{R}}^{(i)}_{t,\text{out}}\right)_{u}\right),$ (23)
again minimising a squared error loss. As in cross-decoding we evaluate $R^{2}$ score (sklearn.metrics.r2_score) and the decoding error $\mathcal{D}_{\bm{r}\rightarrow\bm{\hat{z}}}:=1-(R^{2})_{\bm{r}\rightarrow\bm{\hat{z}}}$ (Fig. 6A).
Summary of Neural Latent Benchmark (NLB) datasets
Here are brief descriptions of the datasets used in this study. All datasets were collected from macaque monkeys performing sensorimotor or cognitive tasks. More comprehensive details can be found in the Neural Latents Benchmark paper [6].
mc_maze [35]
Motor cortex recordings during a delayed reaching task where monkeys navigated around virtual barriers to reach visually cued targets. The task involved 108 unique maze configurations, with several repeated trials for each one, thus serving as a ”neuroscience MNIST”. We choose this dataset to visualise the latents in Fig. 7.
mc_rtt [36]
Motor cortex recordings during naturalistic, continuous reaching toward randomly appearing targets without imposed delays. The task lacks trial structure and includes highly variable movements, emphasizing the need for modeling unpredictable inputs and non-autonomous dynamics.
dmfc_rsg [37]
Recordings from dorsomedial frontal cortex during a time-interval reproduction task, where monkeys estimated and reproduced time intervals between visual cues using eye or hand movements. The task involves internal timing, variable priors, and mixed sensory-motor demands.
area2_bump [38]
Somatosensory cortex recordings during a visually guided reach task in which unexpected mechanical bumps to the limb occurred in half of the trials. The task probes proprioceptive feedback processing and requires modeling input-driven neural responses.
Dimensions of datasets
@@
We analyse several datasets in this work. Three synthetic datasets generated by an RNN, HMM (Methods, Fig. 2) and LGSMM (S4 Fig.) and the four datasets from the Neural Latent Benchmarks (NLB) suite [6, 35, 36, 37, 38]. In table 1, we summarise the dimensions of all these datsets. To evaluate $k$-shot on the existing SOTA methods while maintaining the NLB evaluations, we conserved the forward-prediction aspect. During model training, models output rate predictions for $T^{\text{fp}}$ future time bins in each trial, i.e., equation 1 and equation 2 are evaluated for $1\leq t\leq T^{\text{fp}}$ while input remains as ${\bm{X}}_{1:T,\text{in}}$. Although we do not discuss the forward-prediction metric in our work, we note that the SOTA models receive gradients from this portion of the data.
@@
In all the NLB datasets as well as the RNN dataset we reuse held-out neurons as $k$-out neurons. We do this to preserve NLB evaluation metrics on the SOTA models, as opposed to re-partitioning the dataset resulting in different scores from previous works. This way existing co-smoothing scores are preserved and $k$-shot co-smoothing scores can be directly compared to the original co-smoothing scores. The downside is that we are not testing the few-shot on ‘novel’ neurons. Our numerical results (Fig. 7) show that our concept still applies.
@@
Table 1: Dimensions of real and synthetic datasets. Number of train and test trials $S^{\text{train}}$, $S^{\text{test}}$, time-bins per trial for co-smoothing $T$, and forward-prediction $T^{\text{fp}}$, held-in, held-out and $k$-out neurons $N^{\text{in}}$, $N^{\text{out}}$, $N^{\text{$k$-out}}$. In all the NLB[6] datasets as well the RNN dataset we use the same set of neurons for $N^{\text{out}}$ and $N^{\text{$k$-out}}$.
@@
Synthetic Noisy GRU RNN (Methods) [29] $800$ $200$ $500$ $50$ $10$ $10$
Synthetic HMM (Methods) $2000$ $100$ $10$ 2020 $50$ $50$
Synthetic LGSSM (S4 Fig.) 2020 $500$ $10$ 55 $30$ $30$
mc_maze_20 [35] $1721$ $574$ $35$ $10$ $137$ $45$ $45$
mc_rtt_20 [36] $810$ $270$ $30$ $10$ $98$ $32$ $32$
dmfc_rsg_20 [37] $748$ $258$ $75$ $10$ $40$ $14$ $14$
area2_bump_20 [38] $272$ $92$ $30$ $10$ $49$ $16$ $16$
Theoretical analysis of few shot learning in HMMs.
@@
Consider a student-teacher scenario as in section Good co-smoothing does not guarantee correct latents. We let $T=2$ and use a stationary teacher $z^{(i)}_{1}=z^{(i)}_{2}$. Now consider two examples of inferred students. To ensure a fair comparison, we use two latent states for both students. In the good student, $\xi$, these two states statistically do not depend on time, and therefore it does not have extraneous dynamics. In contrast, the bad student, $\mu$, uses one state for the first time step, and the other for the second time step. A particular example of such students is:
@@
$\hat{B}_{m,n}$ $=\frac{\sum_{i\in\text{$k$-shot trials}}\sum_{t=1}^{T}\mathbb{I}[X^{(i)}_{t,n}=1]\xi^{(i)}_{t,m}}{\sum_{i^{\prime}\in\text{$k$-shot trials}}\sum_{t^{\prime}=1}^{T}\xi^{(i^{\prime})}_{t^{\prime},m}}$ $\forall\ 1\leq m\leq M\text{ and }n\in\text{$k$-out neurons}$ (27)
@@
Fig. 5A shows these analytical results against the left hand side of equation 36 and equation 37 evaluated numerically.
Theoretical analysis of ridgeless least squares regression with extraneous noise.
$x_{i}=z^{*}_{i}+\epsilon_{i},$ (38)
${\bm{z}}_{i}:=\begin{bmatrix}z^{*}_{i}&\bm{\xi}_{i}\end{bmatrix}^{T},$ (39)
$\hat{\bm{w}}=\operatorname*{arg\,min}\left\{\|w\|_{2}:w\text{ minimises }\sum_{i=1}^{k}\|x_{i}-{\bm{w}}^{T}{\bm{z}}_{i}\|^{2}_{2}\right\}.$ (40)
Note that, by construction, the true mapping is:
${\bm{w}}^{*}=\begin{bmatrix}1&0&\dots&0\end{bmatrix}^{T}.$ (41)
Test loss or risk is a mean squared error:
given a test sample ${\bm{z}}_{0}$. The error can be decomposed as:
The scenario described above is a special case of [31]. What follows is a direct application of their theory, which studies the risk $R$, in the limit $k,p\to\infty$ such that $p/k\to\gamma\in(0,\infty)$, to our setting. The alignment of the theory with numerical simulations is demonstrated in Fig. 5B.
Claim 1.
$\gamma<1$, i.e., the underparameterised case $k>p$.
$B=0$ and the risk is just variance and is given by:
with no dependence on $\sigma_{\text{ext}}$.
Proof.
This is a direct restatement of Proposition 2 in [31]. ∎
Claim 2.
$\gamma>1$, i.e., the overparameterised case $k<p$.
The following is true as $k,p\to\infty\text{ and }p/k\to\gamma$:
Proof.
For the non-isotropic case [31] define the following distributions based on the eigendecomposition of $\Sigma$.
$d\widehat{G}(s)$ $=\delta(s-1)$ (48)
[…]
$\gamma c_{0}=\frac{1}{(\gamma-1)\sigma_{\text{ext}}^{2}}$ (49)
We then compute the limiting values of $B$ and $V$:
$V$ $=\sigma_{\text{obs}}^{2}\gamma c_{0}\sigma_{ext}^{2}.$ (51)
Substituting $\gamma c_{0}$ completes the proof. ∎
The extraneous noise, $\sigma_{\text{ext}}$, influences the risk of ridgeless regression only in the regime $k<p$, and its effect is confined to the bias term, leaving the variance unaffected. In contrast, observation noise contributes exclusively to the variance term. Consequently, the dependence of the risk on $\sigma_{\text{ext}}$ persists even in the absence of observation noise, i.e., when $\sigma_{\text{obs}}=0$.
Fig. 5B presents the theoretical predictions alongside the empirical average $k$-shot performance of minimum-norm least-squares regression, computed numerically using the function numpy.linalg.lstsq.
Theoretical analysis of prototype learning for binary classification with extraneous noise.
Teacher latents are distributed as $p(z_{i}^{*})=\frac{1}{2}\delta(z_{i}^{*}-\frac{1}{\sqrt{2}})+\frac{1}{2}\delta(z_{i}^{*}+\frac{1}{\sqrt{2}})$, that is either $\frac{1}{\sqrt{2}}$ or $\frac{1}{\sqrt{2}}$ with probability $\frac{1}{2}$, representing two classes $a$ and $b$ respectively.
This setting is a special case of [53]. They provide a theoretical prediction for average few-shot classification error rate for class a, $\epsilon_{\text{a}}$, given by $\epsilon_{\text{a}}=H(SNR)$ where $H(x)=\frac{1}{\sqrt{2\pi}}\int_{x}^{\infty}dt\exp(-t^{2}/2)$ is a monotonously decreasing function.
In our case this reduces to:
$SNR$ $\approx\frac{\sqrt{Mk}}{\sigma_{\text{ext}}^{2}}$ (54)
The classification error, $\epsilon$, decreases monotonically with the number of samples $k$, tending to zero as $k\to\infty$ for all finite values of $\sigma_{\text{ext}}$. In contrast, $\epsilon$ increases monotonically with extraneous noise $\sigma_{\text{ext}}$, deviating significantly from zero once $\sigma_{\text{ext}}^{2}\approx\sqrt{Mk}$.
Fig. 5C presents the numerically computed error in comparison with the theoretical prediction given in equation 54.
Supplementary information
S1 Fig.
Student-Teacher RNNs: co-smoothing as a function of model size.
Finding the correct model is not just about tuning the latent size hyperparameter. Models over a range of sizes achieve high co-smoothing (Fig. 8).
Fig 8: Performance and model size. Co-smoothing performance $\mathcal{Q}$ as function of the model size for the NODE SAE students with different dimensionalities on the same 64-unit noisy GRU performing 3BFF (Methods). Models of sizes in the range 5-15 yield the highest performance.
S2 Fig.
How to choose $k$ for your dataset?
The main text section on why few-shot works reveals that extraneous models are best discriminated when the shot number, $k$, is small (Fig. 5). So how small can we go? In the case of sparse data like neural spike counts we may obtain $k$-trial subsets in which some neurons are silent. In this scenario the few-shot decoder $g^{\prime}$ receives no signal for those neurons. To avoid this pathological scenario, for each dataset, we pick the smallest possible $k$ that ensures that the probability of encountering silent neurons in a $k$-trial subset is safely near zero. This must be computed for each dataset independently since some datasets are more sparse than others. In Fig. 9 we compute the frequency of such silences for different $k$, for each NLB[6] dataset, and show the values of $k$ chosen for the analysis in the main text.
Fig 9: Choosing $k$ for each of the NLB datasets. Our theoretical analysis suggests taking $k$ to be as small as possible (main text Fig. 3) to maximise the difference in $k$-shot performance. However, in practice, we would like to avoid training on subsets with totally silent neurons. Thus we pick the smallest $k$ that avoids totally silent neurons (dashed lines). This has to be checked for each dataset given its particular distribution of firing rates.
S1 Appendix.
Decoding across HMM latents: fitting and evaluation. Consider two HMMs $u$ and $v$, of sizes $M(u)$ and $M(v)$, both candidate models of a dataset $\mathcal{X}$. Following equation 15, each HMM can be used to infer latents from the data, defining encoder mappings $f^{u}$ and $f^{v}$. These map a single trial $i$ of the data $({\bm{X}}_{:,\text{in}})^{(i)}\in\mathcal{X}$ to $(\bm{\xi}^{(i)}_{t})_{u}$ and $(\bm{\xi}^{(i)}_{t})_{v}$.
$\bm{p}^{(i)}_{t}=h\left(\left(\bm{\xi}^{(i)}_{t}\right)_{u}\right)$ (55)
$h(\xi)=\sigma(W\bm{\xi}+\bm{b})$ (56)
where $D_{KL}$ is implemented with scipy.special.rel_entr.
In section Good co-smoothing does not guarantee correct latents and Fig. 1, the data $X$ is sampled from a single teacher HMM, T, and we evaluate $\mathcal{D}_{\text{T}\rightarrow\text{S}}$ and $\mathcal{D}_{\text{S}\rightarrow\text{T}}$ for each student notated simply as S.
S3 Fig.
Good co-smoothing does not guarantee correct latents in Hidden Markov Models (HMMs). In the main text, we show how good prediction of held-out neural activity, i.e., co-smoothing, does not guarantee a match between model and true latents. We did this in the student-teacher setting of RNNs and NODE SAEs (Fig. 2). Here we replicate the results in HMMs (see Methods) (Fig. 10). The arrows mark the “Good” and “Bad” transition matrices shown in the Fig. 2 (lower).
[…]
S4 Fig.
Student-teacher results in Linear Gaussian State Space Models.
We demonstrate that our results are not unique to the RNN or HMM settings by simulating another simple scenario: linear gaussian state space models (LGSSM), i.e., Kalman Smoothing.
The model is defined by parameters $({\bm{\mu}}_{0},{\bm{\Sigma}}_{0},{\bm{F}},{\bm{G}},{\bm{H}},{\bm{R}})$. A major difference to HMMs is that the latent states $\bm{z}\in\mathbb{R}^{M}$ are continuous. They follow the dynamics given by:
${\bm{z}}_{0}$ $\sim\mathcal{N}({\bm{\mu}}_{0},{\bm{\Sigma}}_{0})$ (58)
${\bm{z}}_{t}$ $\sim\mathcal{N}({\bm{F}}{\bm{z}}_{t-1}+{\bm{b}},{\bm{G}})$ (59)
${\bm{x}}_{t}$ $\sim\mathcal{N}({\bm{H}}{\bm{z}}_{t}+{\bm{c}},{\bm{R}})$ (60)
As with HMMs we use a teacher LGSSM with $M=4$, with parameters chosen randomly (using the dynamax defaults) and then fixed. Student LGSSMs are also initialised randomly and optimised with Adam [50] to minimise negative loglikelihood on the training data (see table 1 for dimensions of the synthetic data set). $\mathcal{D}_{\text{S}\rightarrow\text{T}}$ and $\mathcal{D}_{\text{T}\rightarrow\text{S}}$ is computed with linear regression (sklearn.linear_model.LinearRegression) and predictions are evaluated against the target using $R^{2}$ (sklearn.metrics.r2_score). We define $\mathcal{D}_{u\rightarrow v}:=1-(R^{2})_{u\rightarrow v}$. Few-shot regression from $\bm{z}$ to $\bm{x}^{\text{$k$-out}}$ is also performed using linear regression.
In line with our results with RNNs and HMMs (Fig. 2 and Fig. 4), in Fig. 11 we show that that among the models with high test loglikelihood (>55>-55), $\mathcal{D}_{\text{S}\rightarrow\text{T}}$, but not $\mathcal{D}_{\text{T}\rightarrow\text{S}}$, is highly correlated to test loglikelihood, while $\mathcal{D}_{\text{T}\rightarrow\text{S}}$ shows a close relationship to Average 10 shot MSE error.
Fig 11: Left to right: Student-teacher results for Linear Gaussian State Space Models. We report loglikelihood instead of co-smoothing, and $k$-shot MSE instead of $k$-shot co-smoothing.
S5 Fig.
HMM network visualisations
In the main text Fig. 2 we visualised the teacher and two student HMMs as graphs of fractional traffic volume on states and transitions. For clarity we dropped the low probability edges with values lower than $0.01$. In Fig. 12, we show the same models with all the edges visualised.
Fig 12: HMM visualisations with all transitions included.
S6 Fig.
Few-shot co-smoothing is not simply hard co-smoothing (variations of HMM student-teacher experiments). Few-shot co-smoothing is a more difficult metric than standard co-smoothing. Thus, it might seem that any increase in the difficulty of will yield similar results. To show this is not the case, we use standard co-smoothing with fewer held-in neurons (Fig. 13). The score is lower (because it’s more difficult), but does not discriminate models.
Fig 13: Making co-smoothing harder does not discriminate between models. Top three: Increasing the number of held out neurons from $N^{\text{out}}=50$ to $N^{\text{out}}=100$. First two panels: Same as main text Fig. 1CD. Lower panel: Same as main text Fig. 4B. Bottom three: Decreasing the number of held-in and held-out neurons to $N^{\text{in}}=5$, $N^{\text{out}}=5$, $N^{\text{$k$-out}}=50$. Panels as in top row. The score does decrease because the problem is harder, but co-smoothing is still not indicative of good models while few-shot is.
S2 Appendix.
Code repositories.
The experiments done in this work are largely based on code repositories from previous works. The following repositories were used or developed in this work:
S3 Appendix.
Time cost of computing few-shot co-smoothing
The compute time depends on several factors. It scales with the number of trials $k$, $T$ the number of samples per trial, the number of neurons $N^{\text{$k$-out}}$ and number of repeated resamples $s$ and fitting of the regressor. Each repetition and neuron is an independent regression and therefore can computed in parallel, provided the compute resources are available. For mc_maze_20 each repetition for $k=64$, took $0.62\pm 0.06$ seconds and we iterated over $s=12$ such regressions, requiring a total of $7.44$ seconds.